OpenStudy (anonymous):

A cylindrical tank holds 20,000 gallons of water, which can be drained from the bottom of the tank in 20 minutes. The function gives the volume V of water (in gallons) in the tank at time t in minutes, where time t = 0 corresponds to the instant the tank starts to drain. At time t = 3 minutes, at what rate is the volume of water in the tank changing? (Points : 1) –2200 gallons/minute 1700 gallons/minute 2200 gallons/minute –1800 gallons/minute 1800 gallons/minute

OpenStudy (anonymous):

May I plz have some help?

OpenStudy (anonymous):

u will need a calculator, I jus dont know how to srart it...

OpenStudy (loser66):

do you have the function of V respect to t?

OpenStudy (anonymous):

yes one sec

OpenStudy (anonymous):

V(t)=20,000 (1-(t/20) )squared

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yes

OpenStudy (loser66):

\[ V(t) = 20000(1-\frac{t}{20})^2\] at t =3, V =14450

OpenStudy (anonymous):

at t = 3?

OpenStudy (loser66):

so the volume of water drains out is 20000-14450=5550 5550gallons in 3minutes, that means 5550/3 = 1850/minute which is not one of your options. I don't know why

OpenStudy (loser66):

@amistre64 @ash2326

OpenStudy (loser66):

@mathstudent55

OpenStudy (loser66):

Let wait for those mathematicians, hehehe, sorry for my uselessness

OpenStudy (anonymous):

could u help me with another problem plz?

OpenStudy (anonymous):

Question 4.4. The displacement s (in meters) of a particle moving in a straight line is a function of time t (in seconds). The table below gives the particle's displacement for various times: AP Calculus AB Semester 1 The average velocity of the particle between time t = 1 and time t = 1.8 seconds is: (Points : 1) 11.3536 14.192 12.7534 –11.3536 –14.192

OpenStudy (amistre64):

lets make sure the first one is good

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

A cylindrical tank holds 20,000 gallons of water, which can be drained from the bottom of the tank in 20 minutes. The function gives the volume V of water (in gallons) in the tank at time t in minutes, where time t = 0 corresponds to the instant the tank starts to drain. At time t = 3 minutes, at what rate is the volume of water in the tank changing? (Points : 1) oh, they give you the function .... what is the function?

OpenStudy (anonymous):

V(t)=20,000 (1-(t/20) )squared

OpenStudy (amistre64):

\[V = 20000(1-\frac{t}{20})^2\] take an implicit derivative with respect to time

OpenStudy (anonymous):

how do I do this?

OpenStudy (amistre64):

\[V = k(1-\frac{t}{n})^2\] \[V' = 2k(-\frac 1n)(1-\frac{t}{n})t'\] we know t = 3, and t' = dt/dt = 1 \[V' = 2k(-\frac 1n)(1-\frac{3}{n})\]

OpenStudy (anonymous):

ok, that makes sense, so is this the derivative equation I use?

OpenStudy (amistre64):

without all the big numbers yes

OpenStudy (amistre64):

a rate of change should give you the idea of taking a derivative

OpenStudy (anonymous):

l am a bit stuck

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Could u help me with the second one plz?

OpenStudy (anonymous):

Question 4.4. The displacement s (in meters) of a particle moving in a straight line is a function of time t (in seconds). The table below gives the particle's displacement for various times: AP Calculus AB Semester 1 The average velocity of the particle between time t = 1 and time t = 1.8 seconds is: (Points : 1) 11.3536 14.192 12.7534 –11.3536 –14.192

OpenStudy (amistre64):

an average is the slope of the line between the desired end points

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

given 2 points of reference; the slope between them was well defined back in algebra 2

OpenStudy (anonymous):

The average velocity of the particle between time t = 1 and time t = 1.8 seconds is: 12.7?

OpenStudy (amistre64):

whats the value at t=1? and at t= 1.8?

OpenStudy (anonymous):

13.44, and 2.0864

OpenStudy (amistre64):

well, in order to go from high to low, we have to slow down ... our slope has to be negative for starters

OpenStudy (amistre64):

what is the difference from 13.44 to 2.0864 ?

OpenStudy (anonymous):

11.3536

OpenStudy (anonymous):

wow, that was easy!!!

OpenStudy (amistre64):

yeah, we could have estimated prolly at 11.5 divide that by the .8 seconds it took to get there

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

could u assist me with a few more plz?

OpenStudy (amistre64):

 slope is about -11.5/.8